OpenStudy (anonymous):

Simplify. sqrt144a2b4

OpenStudy (anonymous):

\[\sqrt{144a^2b^4} = \sqrt{144}*\sqrt{a^2}*\sqrt{b^4}\] \[12*a*b^2\]

OpenStudy (anonymous):

A) 12ab B) 12a2b4 C) 12ab2 D) 12a b